OpenStudy (mattduhhelper):

x + 1/3 = 9 What does x represent?

OpenStudy (chiunlee12):

okay, so try subtracting both sides by the same value, so that you can find what x equals

OpenStudy (mattduhhelper):

Subtract both sides? What do you mean?

OpenStudy (chiunlee12):

In this case, you subtract 1/3 from both sides to make x equal the value

OpenStudy (chiunlee12):

so simply, x+ 1/3 = 9 Subtract 1/3 from both sides to cancel out 1/3 so it becomes x+ 1/3 -1/3 = 9-1/3 x+0= 9-1/3

OpenStudy (chiunlee12):

so x+0 = x And you just have to put in x= 9-1/3 and solve

OpenStudy (emogirlatemoooocow):

Simiplfly \[\frac{ 1 }{ 3 }\]

OpenStudy (welshfella):

to calculate 9 - 1/3 , first convert 9 to 1/3's ( 9 = 27/3 so we have 27/3 - 1/3 = 26/3

OpenStudy (emogirlatemoooocow):

@welshfella Isn't it 0 or 6.1?

OpenStudy (welshfella):

how doyou get that?

OpenStudy (emogirlatemoooocow):

26/3 = (27/3-1/3 ) - 26-3 = 0

OpenStudy (welshfella):

thats incorrect

OpenStudy (welshfella):

@MattDuhHelper I have assumed that the 3 in the question divides the 1 only. Thats the way it is written I the 3 divides the x + 1 it should be written (x + 1) / 3 = 9